OpenStudy (anonymous):

Simplify. √121/81 A. 11/81 B. 61/81 C. 11/9 D. 61/41

OpenStudy (danielbarriosr1):

\[\frac{ \sqrt{121} }{81}=\frac{ 11}{81}\]

OpenStudy (anonymous):

 Dont you have to square root the whole fraction. this is how it looks like on my homework..

OpenStudy (danielbarriosr1):

yes, i almost forgot it i was thinking about it, so it is: \[\frac{ 11 }{ 9 }\]

OpenStudy (anonymous):

thanks so much!

OpenStudy (danielbarriosr1):

ok no problem